# Asymptotics for partitions over the Fibonacci numbers and related sequences

### v2

@@ 1. Introduction
We consider two functions each of which describes different types of partitions over Fibonacci numbers, $F_{k}$. In particular, for a positive integer $n$, write
(1) $n=a_{2}F_{2}+a_{3}F_{3}+\cdots+a_{k}F_{k}+\cdots.$
We let $p_{F}(n)$ denote the number of non-distinct partitions of $n$ over the Fibonacci sequence, that is, solutions to (1) in nonnegative integers $a_{k}$. In recent work, Chow and Slattery [2] gave results on the number distinct partitions, $q_{F}(n)$ over the Fibonacci sequence, and noted that their work shows that neither $q_{F}(n)$, nor its partial sums, have a ‘nice’ asymptotic formula. In particular, they showed that there is some oscillation in the partial sums of $q_{F}(n)$, and gave bounds on these oscillations. In a similar vein, very recently, Sana [12] showed that there are also oscillations in the partial sums of the powers $(q_{F}(n))^{m}$ for each $m$. In this paper, with a related motivation, we determine asymptotic formulas for $p_{F}(n)$ and describe the oscillations that occur.
@@
$p_{F}(n)=\psi_{p_{F}}\left(\frac{\log n}{\log\varphi}\right)(\log n)^{-\frac{1}{2}}n^{\left(\frac{\log n}{2\log\varphi}+\frac{\log 5-\log\varphi+4\gamma}{2\log\varphi}-2+O\left(\frac{1}{\log n}\right)\right)}\left(1+O\left(\frac{1}{\log n}\right)\right),$
where $\psi_{p_{F}}(x)$ is an explicit positive $1$-periodic function, $\varphi$ is the golden mean and $\gamma$ is the Euler–Mascheroni constant.
Theorem 1 seems to be the first asymptotic result chracterising non-distinct partitions over Fibonacci numbers. Distinct partitions over Fibonacci numbers have received considerable attention; see [2] and the references therein for details. In very recent work, addressing a question of Chow and Slattery [2], Kempton [5] has shown that $n^{-\frac{\log 2}{\log\varphi}}\sum_{m\leqslant n}q_{F}(m)$ is $\log$-periodic. Our proof of Theorem 1 gives a description of the related $\log$-periodic function for $p_{F}(n)$.
This paper is organised as follows. In Sections 2 and 3, we focus the proof of Theorem 1. In particular, inspired by ideas of de Bruijn [1], we prove an exact formula for the generating function of $p_{F}(n)$ in Section 2. We then use a saddle-point method in Section 3 to give the asymptotic result for $p_{F}(n)$. Finally, in Section 4, we give the complete extension of Theorem 1 to the case of a linear recurrence with dominant root. In particular, suppose that $P_{k}$ is a strictly increasing linearly recurrent sequence of positive integers of degree at least $2$ with $P_{1}=1$, such that the characteristic polynomial $\chi_{P}(x)$ of $P_{k}$ has a single dominant root $\beta>0$ and $P_{k}=\lambda\beta^{k}+\lambda_{2}\beta_{2}^{k}+\cdots+\lambda_{r}\beta_{r}^{k},$ where $\lambda,\lambda_{2},\ldots,\lambda_{r}$ are constants and $\beta_{2},\ldots,\beta_{r}$ are the algebraic conjugates of $\beta$. In Section 4, we prove
@@
$p_{P}(n)=\psi_{p_{P}}\left(\frac{\log n}{\log\beta}\right)(\log n)^{-\frac{1}{2}}n^{\left(\frac{\log n}{2\log\beta}+\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{3}{2}+O\left(\frac{1}{\log n}\right)\right)}\left(1+O\left(\frac{1}{\log n}\right)\right),$
where $\psi_{p_{P}}(x)$ is an explicitly determinable positive $1$-periodic function.
@@ 2. An exact formula for for the generating series of $p_{F}(n)$
as $x\rightarrow 1^{-}$. Note that we are starting with $F_{2}=1$, and not with $F_{1}=1$, since we do not want two representations of $1$. We will necessarily need to consider a Fibonacci Dirichlet series. Navas [8] determined most of the properties we need by considering the analytic continuation of the series $\zeta_{F}(z)=\sum_{k\geqslant 1}F_{k}^{-z}$, but here with $1$ doubly represented. It turns out this is not much of a problem. To deal with this, we merely consider the product
@@
then translate the results back to $F_{2}(x)$. It is convenient to do a change of variables, setting $x=e^{-s}$, so that we are considering the function $F(e^{-s})$, with particular interest in the asymptotics as $s\rightarrow 0^{+}$.
@@
$\log F_{2}(e^{-s})=\frac{(\log s)^{2}}{2\log\varphi}-(\log s)\left(\frac{c_{3}}{\log\varphi}-1\right)+c_{2}+2\gamma+f(s)+O(s).$
@@
where $c_{1}:=\sum_{k\geqslant 1}\frac{(-1)^{k}}{k(\varphi^{2k}+(-1)^{k+1})}\approx-0.20436188.$ Thus, using (9),
@@
$\underset{{{z}={0}}}{\textnormal{Res}}\left\{s^{-z}\Gamma(z)\zeta(1+z)\zeta_{F}(z)\right\}=\frac{1}{\log\varphi}\left(\frac{(\gamma-\log s)(3\gamma-\log s)}{2}-\gamma_{1}+\frac{\pi^{2}}{12}\right)\\ +\frac{\log 5-\log\varphi}{2\log\varphi}(2\gamma-\log s)+\left(\frac{\log\varphi-3\log 5}{12}+\frac{(\log 5)^{2}}{8\log\varphi}+c_{1}\right)$
where $c_{1}:=\sum_{k\geqslant 1}\frac{(-1)^{k}}{k(\varphi^{2k}+(-1)^{k+1})}\approx-0.20436188.$ Gathering powers of $\log s$ finishes the proof. ∎
@@
For the calculation of $\zeta_{F}(-1)$ see Navas [8, Eq. 9]
@@
$\zeta_{F}(z)$ $=5^{z/2}\sum_{k=0}^{\infty}\binom{-z}{k}\frac{(-1)^{k}}{\phi^{z+2k}+(-1)^{k+1}}=5^{z/2}\sum_{k=0}^{\infty}\frac{\Gamma(1-z)}{\Gamma(1-z-k)k!}\cdot\frac{(-1)^{k}}{\phi^{z+2k}+(-1)^{k+1}}$
$=5^{z/2}\frac{\Gamma(1-z)}{\Gamma(1-z-2n)(2n)!}\cdot\frac{1}{\phi^{z+4n}-1}+5^{z/2}\sum_{\begin{subarray}{c}k=0\\ k\neq 2n\end{subarray}}^{\infty}\frac{\Gamma(1-z)}{\Gamma(1-z-k)k!}\cdot\frac{(-1)^{k}}{\phi^{z+2k}+(-1)^{k+1}}.$
@@
$\lim_{z\rightarrow-4n}\frac{d}{dz}\left\{\frac{z+4n}{\phi^{z+4n}-1}\right\}=-\frac{1}{2},$
@@
$b_{-4n}:=\underset{{{z}={-4n}}}{\textnormal{Res}}\left\{\zeta_{F}(z)\right\}=\frac{5^{-2n}}{\log\phi}\binom{4n}{2n}=\frac{5^{-2n}}{\log\phi}\cdot\frac{(4n)!}{((2n)!)^{2}},$
@@
and $\zeta(1+z)=\zeta(1-4n)+\zeta^{\prime}(1-4n)(z+4n)+O((z+4n)^{2})$ along with the fact that $\zeta(1-4n)=B_{2n}/4n$; see Titchmarch [14, p. 19].
@@
Using the fact that $|B_{4n}|\sim\tfrac{2(4n)!}{(2\pi)^{4n}}$ with Stirling’s approximation of the factorial, we have that
$\frac{b_{-4n}}{(4n)!}\cdot\zeta^{\prime}(1-4n)=O\left(\frac{(4n)!\log n}{(2\pi\sqrt{5})^{4n}n((2n)!)^{2}}\right)=O\left(\frac{\log n}{(2\pi\sqrt{5})^{4n}n^{3/2}}\right)=O(1).$
@@
$|c_{-4n}|=\left|\frac{1}{5^{2n}}\sum_{\begin{subarray}{c}k=0\\ k\neq 2n\end{subarray}}^{\infty}{4n\choose k}\cdot\frac{(-1)^{k}}{\phi^{-4n+2k}+(-1)^{k+1}}\right|=\frac{1}{5^{2n}}\sum_{\begin{subarray}{c}k=0\\ k\neq 2n\end{subarray}}^{4n}{4n\choose k}\cdot\varphi^{4n-2k}\left|\frac{1}{1-\phi^{4n-2k}}\right|$
since (${4n\choose k}=0$ for $k>4n$. Now the value of $\left|\tfrac{1}{1-\phi^{4n-2k}}\right|$ is maximal when $k=2n-1$ (recall, $k\neq 2n$), and there, it is approximately $2.6180$, which yields
$|c_{-4n}|<\frac{3}{5^{2n}}\sum_{\begin{subarray}{c}k=0\\ k\neq 2n\end{subarray}}^{4n}{4n\choose k}\cdot\varphi^{4n-2k}\leqslant\frac{3}{5^{2n}}\sum_{\begin{subarray}{c}k=0\end{subarray}}^{4n}{4n\choose k}\cdot\phi^{4n-k}\left(\frac{1}{\varphi}\right)^{k}=\frac{3}{5^{2n}}\left(\varphi+\frac{1}{\varphi}\right)^{4n}=3,$
@@
$|\alpha_{n}|$ $<\frac{|B_{2n}|}{4n\cdot(4n)!}\left(\frac{5^{-2n}}{\log\phi}\cdot\frac{(4n)!}{((2n)!)^{2}}\left(\sum_{k=1}^{4n}\frac{1}{k}-\gamma\right)+3+\frac{5^{-2n}}{2}\frac{(4n)!}{((2n)!)^{2}}\right)+O(1)$
$=\zeta(2n)\frac{2(2n)!}{4n\cdot(2\pi)^{2n}\cdot(4n)!}\left(\frac{5^{-2n}}{\log\phi}\cdot\frac{(4n)!}{((2n)!)^{2}}\left(\sum_{k=1}^{4n}\frac{1}{k}-\gamma\right)+3+\frac{5^{-2n}}{2}\frac{(4n)!}{((2n)!)^{2}}\right)+O(1)$
$\sim\frac{2(2n)!}{4n\cdot(2\pi)^{2n}\cdot(4n)!}\left(\frac{5^{-2n}}{\log\phi}\cdot\frac{(4n)!}{((2n)!)^{2}}\log(4n)+3+\frac{5^{-2n}}{2}\frac{(4n)!}{((2n)!)^{2}}\right)+O(1)$
$=\frac{2}{4n\cdot(2\pi)^{2n}}\left(\frac{5^{-2n}}{\log\phi}\cdot\frac{\log(4n)}{(2n)!}+\frac{3\cdot(2n)!}{(4n)!}+\frac{5^{-2n}}{2(2n)!}\right)+O(1)$
$\sim\frac{2}{4n\cdot(2\pi)^{2n}}\cdot\frac{5^{-2n}}{\log\phi}\cdot\frac{\log(4n)}{(2n)!}+O(1)\sim\frac{1}{4\sqrt{\pi}\log{\varphi}}\cdot\frac{\log n}{\sqrt{n}}\cdot\left(\frac{e}{20\pi n}\right)^{2n}+O(1)=O(1).$
@@
To see that $g(s)=O(s^{2})$ as $s\to 0^{+}$, we note that, using L’Hôpital’s rule, we have $\lim_{s\to 0^{+}}s^{2}\log s=0$. Using this, along with an application of the first part of the proposition,
@@
We first consider the $k=0$ term of $h(s)$, which we denote by $[k=0]h(s)$. By a well-known properties of $\Gamma$ and $\zeta$, we have, for any real nonzero $t$, that, as $t\to\infty$, both $|\Gamma(it)|^{2}\sim\frac{\pi}{te^{\pi t}}$ and $\zeta(1+it)=O(\log|t|).$ Noting that for any complex number $z$, (${z\choose 0}=1$, we thus have that
$\big{|}[k=0]h(s)\big{|}\leqslant\sum_{n=1}^{\infty}\frac{O(\log n)}{\sqrt{n}\cdot e^{\frac{\pi^{2}n}{\log\varphi}}}=O(1).$
@@
(12) $<\frac{d}{\big{(}e^{\frac{\pi^{2}}{2\log\varphi}}-1\big{)}}\left(\frac{s^{2}}{10\pi\cdot\log\varphi\cdot\big{(}e^{\frac{\pi^{2}}{2\log\varphi}}-1\big{)}}\right)^{k},$
where, for $k\geqslant 1$, we have used that $\sum_{n\geqslant 0}n^{k}x^{n}={x\,A_{k}(x)}{(1-x)^{-(k+1)}},$ where $A_{k}(x)\in\mathbb{Z}_{>0}[x]$ is the $k$-th Eulerian Polynomial, which satisfies $\deg A_{k}(x)=k-1$ and $A_{k}(1)=k!$.
Finally, to see that $h(s)=h(\varphi s)+O(s^{2})$ as $s\to 0^{+}$, we note that by (12), the terms $[k\neq 0]h(s)$ contribute $O(s^{2})$ collectively, and for $[k=0]h(s)$, we note that
$=\cos\left(\frac{2\pi n}{\log\varphi}\log\left(\frac{\alpha}{\sqrt{5}}\right)\right)+i\sin\left(\frac{2\pi n}{\log\varphi}\log\left(\frac{\alpha}{\sqrt{5}}\right)\right)=\left(\frac{s}{\sqrt{5}}\right)^{\frac{i\pi(2n)}{\log\varphi}}.\qed$
@@
where, for now, $a>0$. We use the expansions of $\Gamma(z)$ and $\zeta(1+z)$ from above, to get that, as $s\to 0^{+}$,
$-\log(1-e^{-s})=-\log s+2\gamma+O(s)$
which, when multiplied by $-1$, yields the desired result. ∎

### v3

@@ 1. Introduction
We consider the number $p_{F}(n)$ of non-distinct partitions of $n$ over the Fibonacci sequence, $F_{k}$. Specifically, for a positive integer $n$, $p_{F}(n)$ is the number of solutions of
(1) $n=a_{2}F_{2}+a_{3}F_{3}+\cdots+a_{k}F_{k}+\cdots,$
in nonnegative integers $a_{k}$. In recent work, Chow and Slattery [2] gave results on the number distinct partitions, $q_{F}(n)$ over the Fibonacci sequence, and noted that their work shows that neither $q_{F}(n)$, nor its partial sums, have a ‘nice’ asymptotic formula. In particular, they showed that there is some oscillation in the partial sums of $q_{F}(n)$, and they gave bounds on these oscillations. In a similar vein, very recently, Sana [14] showed that there are also oscillations in the partial sums of the powers $(q_{F}(n))^{m}$ for each $m$. In this paper, with a related motivation, we determine asymptotic formulas for $p_{F}(n)$ and describe the oscillations that occur.
@@
$\log p_{F}(n)\sim\frac{(\log n)^{2}}{2\log\varphi}.$
In particular,
$p_{F}(n)=A_{F}\left(\frac{\log n}{\log\varphi}\right)n^{B_{F}(n)}(\log n)^{C_{F}(n)}\left(1+O\left(\frac{(\log\log n)^{2}}{\log n}\right)\right),$
where
$A_{F}(x)$ $:=\sqrt{\frac{1}{2\pi\log\varphi}}\cdot\exp\Bigg{(}\frac{(\log\log\varphi)^{2}}{2\log\varphi}+\log\log\varphi\left(\frac{c_{3}}{\log\varphi}-1\right)+c_{2}+2\gamma+\psi_{1}(x)\Bigg{)},$
$B_{F}(n)$ $:=\frac{1}{2\log\varphi}\left(\log n-2\log\log n-\frac{4\psi_{0}\left(\frac{\log n}{\log\varphi}\right)}{\log\varphi}+2\log\log\varphi+2c_{3}-4\log\varphi+2\right),$
$C_{F}(n)$ $:=\frac{1}{2\log\varphi}\left(\log\log n-\frac{4\psi_{0}\left(\frac{\log n}{\log\varphi}\right)}{\log\varphi}+2\log\log\varphi-2c_{3}+3\log\varphi+4\right),$
$c_{2}$ $:=\frac{3\gamma^{2}}{2\log\varphi}-\frac{\gamma_{1}}{\log\varphi}+\frac{\pi^{2}}{12\log\varphi}+2\gamma\left(\frac{\log 5-\log\varphi}{2\log\varphi}\right)+\sum_{k\geqslant 1}\frac{(-1)^{k}}{k(\varphi^{2k}+(-1)^{k+1})},\mbox{ and}$
$c_{3}$ $:=\frac{1}{2}\log 5-\frac{1}{2}\log\varphi+2\gamma,$
where $\varphi$ is the golden mean, $\gamma$ is the Euler–Mascheroni constant, $\gamma_{1}$ is the first Stieltjes constant, and $\psi_{0}(x)$ and $\psi_{1}(x)$ are explicitly computable $1$-periodic functions.
Theorem 1 seems to be the first asymptotic result chracterising non-distinct partitions over Fibonacci numbers. Distinct partitions over Fibonacci numbers have received considerable attention; see [2] and the references therein for details. In very recent work, addressing a question of Chow and Slattery [2], Kempton [7] has shown that $n^{-\frac{\log 2}{\log\varphi}}\sum_{m\leqslant n}q_{F}(m)$ is $\log$-periodic. Our proof of Theorem 1 gives a way to describe the related $\log$-periodic function for $p_{F}(n)$.
Our main result on Fibonacci partitions followa from a result of Coons and Kirsten [3], which uses a saddle-point method, which itself was inspired by the work of Nanda [9] and Richmond [12, 13]. In particular, Richmond [12] was not only able to give a new proof of Hardy and Ramanujan’s above-mentioned result, he gave asymptotics for all of the moments of $p(n)$. The method therein, and herein, relies on the existence of invertible asymptotics for the related saddle point. In our situation, the lead order asymptotic of the saddle point is monotonic, so we have invertibility, but, additionally, we are able compute the second-order term, which is oscillatory. These terms are asymptotically close enough, so that they both contribute to the outcome of Theorem 1. We note that our results are heavily related to analytic properties of the Fibonacci zeta function (defined and discussed in more detail below); in particular, the Fibonacci zeta function is defined by a Dirichlet series which converges in the positive right half-plane. This presents added difficulties compared to the more fully-examine case of partitions related to Dirichlet series whose abscissa of convergence is strictly positive—for an interesting study on the asymptotics of partitions related to to Dirichlet series whose abscissa of convergence is strictly positive, see Debruyne and Tenenbaum [4].
This paper is organised as follows. In Sections 2 and 3, we focus the proof of Theorem 1. In particular, inspired by ideas of de Bruijn [1], we prove an exact formula for the generating function of $p_{F}(n)$ in Section 2. We then use a saddle-point method in Section 3 to give the asymptotic result for $p_{F}(n)$. Finally, in Section 4, we give the complete extension of Theorem 1 to the case of a linear recurrence with dominant root. In particular, suppose that $P_{k}$ is a strictly increasing linearly recurrent sequence of positive integers of degree at least $2$ with $P_{1}=1$, such that the characteristic polynomial $\chi_{P}(x)$ of $P_{k}$ has a single dominant root $\beta>0$ and $P_{k}=\lambda\beta^{k}+\lambda_{2}\beta_{2}^{k}+\cdots+\lambda_{r}\beta_{r}^{k},$ where $\lambda,\lambda_{2},\ldots,\lambda_{r}$ are constants and $\beta_{2},\ldots,\beta_{r}$ are the algebraic conjugates of $\beta$, then we have
@@
$p_{P}(n)=A_{P}\left(\frac{\log n}{\log\beta}\right)n^{B_{P}(n)}(\log n)^{C_{P}(n)}\left(1+O\left(\frac{(\log\log n)^{2}}{\log n}\right)\right),$
where $A_{P}(x)$ is the a positive $1$-periodic function satisfying
$A_{P}(x)$ $:=\sqrt{\frac{1}{2\pi\log\beta}}\cdot\exp\Bigg{(}\frac{(\log\log\beta)^{2}}{2\log\beta}+(2\gamma-\log\lambda)\frac{\log\log\beta}{\log\beta}-\frac{1}{2}\log\log\beta+C_{2}+\psi_{4}(x)\Bigg{)},$
$B_{P}(n)$ $:=\frac{1}{2\log\beta}\left(\log n-2\log\log n-\frac{4\psi_{3}\left(\frac{\log n}{\log\beta}\right)}{\log\beta}+2\log\log\beta+4\gamma-2\log\lambda-2\log\beta+2\right),$
$C_{P}(n)$ $:=\frac{1}{2\log\beta}\left(\log\log n+2-\frac{4\psi_{3}\left(\frac{\log n}{\log\beta}\right)}{\log\beta}+2\log\log\beta-4\gamma+2\log\lambda+2\log\beta+2\right),$
where $\psi_{3}(x)$ and $\psi_{4}(x)$ are explicitly computable $1$-periodic functions, and $C_{2}$ is the constant defined in Proposition 9.
@@ 2. An exact formula for for the generating series of $p_{F}(n)$
as $x\rightarrow 1^{-}$. Note that we are starting with $F_{2}=1$, and not with $F_{1}=1$, since we wish to avoid having two representations of $1$ in our partitions. We will necessarily need to consider a Fibonacci Dirichlet series. Navas [10] determined most of the properties we need by considering the analytic continuation of the series $\zeta_{F}(z)=\sum_{k\geqslant 1}F_{k}^{-z}$, but here with $1$ doubly represented. It turns out this is not much of a problem. To deal with this, we merely consider the product
@@
then translate the results back to $F_{2}(x)$. It is convenient to change variables, setting $x=e^{-s}$, so that we are considering the function $F(e^{-s})$, with particular interest in the asymptotics as $s\rightarrow 0^{+}$.
@@
$\log F_{2}(e^{-s})=\frac{(\log s)^{2}}{2\log\varphi}-(\log s)\left(\frac{c_{3}}{\log\varphi}-1\right)+c_{2}+2\gamma+f(s)+O(s^{2}).$
@@
where $c_{1}$ is defined as in the statement of the lemma. Thus, using (9),
@@
$\underset{{{z}={0}}}{\textnormal{Res}}\left\{s^{-z}\Gamma(z)\zeta(1+z)\zeta_{F}(z)\right\}=\frac{1}{\log\varphi}\left(\frac{(\gamma-\log s)(3\gamma-\log s)}{2}-\gamma_{1}+\frac{\pi^{2}}{12}\right)\\ +\frac{\log 5-\log\varphi}{2\log\varphi}(2\gamma-\log s)+\left(\frac{\log\varphi-3\log 5}{12}+\frac{(\log 5)^{2}}{8\log\varphi}+c_{1}\right).$
Gathering powers of $\log s$ finishes the proof. ∎
@@
For the calculation of $\zeta_{F}(-1)$ see Navas [10, Eq. 9]
@@
$\zeta_{F}(z)$ $=5^{z/2}\sum_{k=0}^{\infty}\binom{-z}{k}\frac{(-1)^{k}}{\varphi^{z+2k}+(-1)^{k+1}}=5^{z/2}\sum_{k=0}^{\infty}\frac{\Gamma(1-z)}{\Gamma(1-z-k)k!}\cdot\frac{(-1)^{k}}{\varphi^{z+2k}+(-1)^{k+1}}$
$=5^{z/2}\frac{\Gamma(1-z)}{\Gamma(1-z-2n)(2n)!}\cdot\frac{1}{\varphi^{z+4n}-1}+5^{z/2}\sum_{\begin{subarray}{c}k=0\\ k\neq 2n\end{subarray}}^{\infty}\frac{\Gamma(1-z)}{\Gamma(1-z-k)k!}\cdot\frac{(-1)^{k}}{\varphi^{z+2k}+(-1)^{k+1}}.$
@@
$\lim_{z\rightarrow-4n}\frac{d}{dz}\left\{\frac{z+4n}{\varphi^{z+4n}-1}\right\}=-\frac{1}{2},$
@@
$b_{-4n}:=\underset{{{z}={-4n}}}{\textnormal{Res}}\left\{\zeta_{F}(z)\right\}=\frac{5^{-2n}}{\log\varphi}\binom{4n}{2n}=\frac{5^{-2n}}{\log\varphi}\cdot\frac{(4n)!}{((2n)!)^{2}},$
@@
and $\zeta(1+z)=\zeta(1-4n)+\zeta^{\prime}(1-4n)(z+4n)+O((z+4n)^{2})$ along with the fact that $\zeta(1-4n)=B_{2n}/4n$; see Titchmarch [16, p. 19].
@@
Using the facts that $\zeta^{\prime}(4n)\sim-2^{-4n}\log 2$ and $|B_{4n}|\sim\tfrac{2(4n)!}{(2\pi)^{4n}}$ with Stirling’s approximation of the factorial, we have that
$\frac{b_{-4n}}{(4n)!}\cdot\zeta^{\prime}(1-4n)=O\left(\frac{(4n)!}{(40\pi)^{2n}n((2n)!)^{2}}\right)=O\left(\frac{1}{(10\pi)^{2n}n^{3/2}}\right)=O(1).$
@@
$|c_{-4n}|=\left|\frac{1}{5^{2n}}\sum_{\begin{subarray}{c}k=0\\ k\neq 2n\end{subarray}}^{\infty}{4n\choose k}\cdot\frac{(-1)^{k}}{\varphi^{-4n+2k}+(-1)^{k+1}}\right|=\frac{1}{5^{2n}}\sum_{\begin{subarray}{c}k=0\\ k\neq 2n\end{subarray}}^{4n}{4n\choose k}\cdot\varphi^{4n-2k}\left|\frac{1}{1-\varphi^{4n-2k}}\right|$
since (${4n\choose k}=0$ for $k>4n$. Now the value of $\left|\tfrac{1}{1-\varphi^{4n-2k}}\right|$ is maximal when $k=2n-1$ (recall, $k\neq 2n$), and there, it is approximately $2.6180$, which yields
$|c_{-4n}|<\frac{3}{5^{2n}}\sum_{\begin{subarray}{c}k=0\\ k\neq 2n\end{subarray}}^{4n}{4n\choose k}\cdot\varphi^{4n-2k}\leqslant\frac{3}{5^{2n}}\sum_{\begin{subarray}{c}k=0\end{subarray}}^{4n}{4n\choose k}\cdot\varphi^{4n-k}\left(\frac{1}{\varphi}\right)^{k}=\frac{3}{5^{2n}}\left(\varphi+\frac{1}{\varphi}\right)^{4n}=3,$
@@
$|\alpha_{n}|$ $<\frac{|B_{2n}|}{4n\cdot(4n)!}\left(\frac{5^{-2n}}{\log\varphi}\cdot\frac{(4n)!}{((2n)!)^{2}}\left(\sum_{k=1}^{4n}\frac{1}{k}-\gamma\right)+3+\frac{5^{-2n}}{2}\frac{(4n)!}{((2n)!)^{2}}\right)+O(1)$
$=\zeta(2n)\frac{2(2n)!}{4n\cdot(2\pi)^{2n}\cdot(4n)!}\left(\frac{5^{-2n}}{\log\varphi}\cdot\frac{(4n)!}{((2n)!)^{2}}\left(\sum_{k=1}^{4n}\frac{1}{k}-\gamma\right)+3+\frac{5^{-2n}}{2}\frac{(4n)!}{((2n)!)^{2}}\right)+O(1)$
$\sim\frac{2(2n)!}{4n\cdot(2\pi)^{2n}\cdot(4n)!}\left(\frac{5^{-2n}}{\log\varphi}\cdot\frac{(4n)!}{((2n)!)^{2}}\log(4n)+3+\frac{5^{-2n}}{2}\frac{(4n)!}{((2n)!)^{2}}\right)+O(1)$
$=\frac{2}{4n\cdot(2\pi)^{2n}}\left(\frac{5^{-2n}}{\log\varphi}\cdot\frac{\log(4n)}{(2n)!}+\frac{3\cdot(2n)!}{(4n)!}+\frac{5^{-2n}}{2(2n)!}\right)+O(1)$
$\sim\frac{2}{4n\cdot(2\pi)^{2n}}\cdot\frac{5^{-2n}}{\log\varphi}\cdot\frac{\log(4n)}{(2n)!}+O(1)\sim\frac{1}{4\sqrt{\pi}\log{\varphi}}\cdot\frac{\log n}{\sqrt{n}}\cdot\left(\frac{e}{20\pi n}\right)^{2n}+O(1)=O(1).$
@@
To see $g(s)=O(s^{2})$ as $s\to 0^{+}$, we note that, using L’Hôpital’s rule, we have $\lim_{s\to 0^{+}}s^{2}\log s=0$. Using this, along with an application of the first part of the proposition,
@@
We first consider the $k=0$ term of $h(s)$, which we denote by $[k=0]h(s)$. Noting that
$=\cos\left(\frac{2\pi n}{\log\varphi}\log\left(\frac{\alpha}{\sqrt{5}}\right)\right)+i\sin\left(\frac{2\pi n}{\log\varphi}\log\left(\frac{\alpha}{\sqrt{5}}\right)\right)=\left(\frac{s}{\sqrt{5}}\right)^{\frac{i\pi(2n)}{\log\varphi}},$
we have that $[k=0]h(\varphi s)=[k=0]h(s)$.
@@
$<\frac{d}{\big{(}e^{\frac{\pi^{2}}{2\log\varphi}}-1\big{)}}\left(\frac{s^{2}}{10\pi\cdot\log\varphi\cdot\big{(}e^{\frac{\pi^{2}}{2\log\varphi}}-1\big{)}}\right)^{k},$
where, for $k\geqslant 1$, we have used that $\sum_{n\geqslant 0}n^{k}x^{n}={x\,A_{k}(x)}{(1-x)^{-(k+1)}},$ where $A_{k}(x)\in\mathbb{Z}_{>0}[x]$ is the $k$-th Eulerian Polynomial, which satisfies $\deg A_{k}(x)=k-1$ and $A_{k}(1)=k!$. Hence, the terms $[k\neq 0]h(s)$ contribute $O(s^{2})$ collectively, and the lemma follows.∎
@@
where, for now, $a>0$. We use the expansions of $\Gamma(z)$ and $\zeta(1+z)$ around $z=0$ from above, with the fact that the integrand has a simple pole at $z=-1$ coming from $\Gamma(z)$ to get that, as $s\to 0^{+}$,
$-\log(1-e^{-s})=-\log s+2\gamma-s\zeta(0)+O(s^{2})$
which, since $\zeta(0)=-1/2$, when multiplied by $-1$, yields the desired result. ∎